# Enhanced Distributed Variational Quantum Eigensolver for Large-Scale MaxCut Problem

### v1

@@ Abstract
MaxCut is a canonical NP-hard combinatorial optimization problem in graph theory with broad applications ranging from physics to bioinformatics. Although variational quantum algorithms offer promising new approaches that may eventually outperform classical schemes, they suffer from resource constraints and trainability issues such as barren plateaus, making large-scale instances intractable on noisy intermediate-scale quantum devices. In this paper, we propose an enhanced distributed variational quantum eigensolver for large-scale MaxCut problems, which extends our prior distributed variational quantum eigensolver framework by integrating a novel hybrid classical-quantum perturbation strategy, enhances optimization scalability and efficiency. Our algorithm solves weighted MaxCut instances with up to 1000 vertices using only 10 qubits, and numerical results indicate that it consistently outperforms the Goemans–Williamson algorithm. We further employ a warm-start initialization strategy, seeding the algorithm with high-quality solutions from the Goemans–Williamson algorithm, with results confirming that the optimal classical solution can be effectively further improved. The practical utility of the proposed algorithm is further validated through its application to haplotype phasing on genome sequencing data of the human ABCA1 gene, producing high-quality haplotypes that rival those obtained by the Goemans-Williamson algorithm with projections. These results establish the proposed algorithm as a scalable, NISQ-compatible framework for near-term quantum-enhanced large-scale combinatorial optimization.
@@ 1 Introduction
For large-scale datasets, MaxCut becomes computationally intractable for classical heuristics, despite their effectiveness on smaller instances. This fundamental barrier drives the search for new computational paradigms, particularly quantum computing, which holds promise for outperforming classical methods on such hard problem. In the current noisy intermediate-scale quantum (NISQ) era[10], variational quantum algorithms (VQAs) [11, 12, 13, 14] represent the leading hybrid quantum-classical framework for tackling combinatorial optimization tasks. By employing a parameterized quantum circuit, also named ansatz, whose parameters are iteratively optimized by a classical optimizer to minimize or maximize a problem-encoded cost function, VQAs strike an attractive balance between expressiveness and hardware compatibility. Such as the variational quantum eigensolver (VQE)[15], originally developed for finding molecular ground states but readily adapted to Ising-type objectives, and the quantum approximate optimization algorithm (QAOA) [16], specifically designed for discrete combinatorial problems such as MaxCut. Both approaches are relatively resilient to noise when operated at shallow circuit depths, making them among the most viable candidates for near-term quantum advantage.
@@
However, such distributed frameworks are also plagued by the barren plateaus, a pervasive challenge in variational quantum algorithms, where the cost function’s gradient vanishes exponentially as the number of qubits increases [24, 25, 26, 27]. This decay severely impedes the classical optimization loop, often trapping it in suboptimal local minima and limiting scalability. The issues of vanishing gradients and local optima are likewise prevalent in classical optimization. In metaheuristic algorithms, a common and effective countermeasure is the deliberate introduction of perturbations. For instance, simulated annealing employs temperature-controlled stochastic acceptance [28], genetic algorithms utilize mutation operators [29], and variable neighborhood search relies on systematic neighborhood shifts [30, 31]. These strategies have all proven effective on challenging problems such as MaxCut and the traveling salesman problem (TSP) [32, 33]. In variational quantum algorithms, an analogous and highly effective perturbation-like technique is warm-starting. It leverages a high-quality classical solution to prepare a biased initial state, thereby accelerating convergence, facilitating escape from local minima, and increasing robustness against barren plateaus. Numerical experiments on small-scale graph instances further confirm that the warm-start method consistently outperforms the standard Goemans–Williamson algorithm[34, 35].
@@
A hybrid classical-quantum perturbation method rooted in the MaxCut problem is proposed and integrated with the DVQE algorithm, enabling the solution of large-scale instances using merely 10 qubits, which significantly reducing hardware requirements while improving algorithmic performance.
@@
Simulation results reveal that the warm-start methodology consistently outperforms the Goemans–Williamson algorithm.
@@ 2.2 Haplotype phasing
Traditionally, haplotype information has been inferred statistically from population genotype data using linkage disequilibrium, as implemented in algorithms like PHASE[36]. However, the switch error rate is relatively high (approximately 5.4%) in unrelated individuals[37]. With the advancement of high-throughput sequencing technologies, particularly the emergence of long fragments and paired-end sequencing, direct haplotype phasing from sequencing data of a single individual has become feasible. The HapCUT algorithm [38] proposed by Bansal and Bafna (2008) represents an efficient and accurate combinatorial optimization method that transforms the problem into iteratively solving MaxCut on associated graphs, minimizing the minimum error correction (MEC) score. On the HuRef personal genome data, it achieved assembly accuracy significantly superior to greedy heuristics (20–25% lower MEC scores) and previous methods, reducing the switch error rate to 1.1–1.4%. This advancement provides more reliable complete haplotype information for personalized genomics, disease association studies, and precision medicine.
@@ 3.1 Framework of Enhanced Distributed Variational Quantum Eigensolver
Classical Neighborhood Search: Starting from apply a classical neighborhood search algorithm (e.g., one based on single-vertex flips) to perform a local exploration, yielding an improved solution .
@@ 3.2 Distributed Variational Quantum Eigensolver
In practice, the Hamiltonian for binary optimization problems is typically expressed as a linear combination of terms:o
@@
Moreover, each subsystem state is prepared and optimized on smaller quantum processors, while the global energy expectation is computed through the factorized expression:
@@ 3.3 Hybrid Classical-Quantum Perturbation
where yields the identity operation , and results in , performing a SWAP operation up to a global phase. The parameter for each gate is initialized from a uniform distribution to ensure the system starts from a state that is only slightly perturbed from the CNS-1 solution. The subsequent variational optimization of these parameters enables probabilistic swapping of partition assignments, thereby facilitating exploration of more diverse solution spaces.
@@ 4.2 Performance on Weighted Complete Graphs
The performance of both algorithms, in terms of the normalized average cut value , is summarized in Figure 2 across a range of vertex counts. The results clearly demonstrate the superiority of the EDVQE algorithm over the standard GW algorithm under various projection configurations. Specifically, for problem sizes ranging from 100 to 1000 vertices, the EDVQE without hybrid classical-quantum perturbation, which represents the initial solution, consistently outperforms the standard GW algorithm employing a single projection. This observation demonstrates the inherent advantage of the base DVQE framework. Moreover, integrating hybrid classical-quantum perturbation within the EDVQE framework yields supplementary performance enhancements. To establish a comprehensive benchmark, we enhance the GW algorithm by increasing its number of projections, which correspondingly elevated its solution quality. As the result, our EDVQE approach maintains consistent superiority over the GW variant with 10 projections and, notably, outperforms the GW algorithm with 100 projections across instances ranging from 600 to 1000 vertices.
To further analyze the performance gains within the EDVQE framework, we break down the contributions of its three key components, which include the initial solution, the classical single-vertex neighborhood search (CNS-1), and the quantum two-vertex perturbation (QP-2). Table 1 summarizes the normalized average cut values of the initial solution, along with the absolute improvements from CNS-1 (CNS-1) and QP-2 (QP-2), together with their corresponding percentage enhancements. The results indicate that CNS-1 provides a consistent yet modest absolute improvement, with an average of 0.0023 (a relative gain of 0.088%). In contrast, QP-2 contributes a more substantial enhancement, delivering an average of 0.0043 (a relative gain of 0.150%)—approximately double that of the classical component. These findings demonstrate that both post-processing modules consistently improve upon the initial solution, with the quantum-based perturbation yielding notably greater gains.
| Value | % Improvement | |||||
@@ 4.3 Performance on Weighted Cluster Graphs
Table 2 presents a component-wise analysis of the performance improvements in the EDVQE. The CNS-1 yields minimal improvement, with an absolute gain of 0.0051, corresponding to a relative gain of 0.085%. In contrast, the QP-2 contributes a substantially larger enhancement, attaining an absolute of 0.1386 and a relative gain of 2.168% relatively, exceeding the classical component by more than 25 times. This pronounced performance gap underscores the quantum module’s distinctive capability to identify cross-community optimal configurations via coordinated multi-qubit operations, thereby effectively addressing the limitations of classical local search in modular graph topologies.
| Value | % Improvement | |||||
@@ 4.5 Performance for Haplotype phasing
| GW () | EDVQE | ||||
| Best value found | avg_cut_value | Success_rate | avg_arrpoximates_ratio | ||
@@
To assess phasing accuracy, the performance of EDVQE is systematically compared against the gold-standard small variant benchmark dataset derived from the Genome in a Bottle Consortium. As shown in Table 4, EDVQE achieves 100% phasing completeness with zero swith error rate and Hamming error rate, with independent validation using the bioinformatics tool WhatsHap confirming identical results. The comparative analysis demonstrate that EDVQE is viable for real-world haplotyping applications, accurately phasing heterozygous SNP variants into fully resolved, error-free haplotypes comparable to established bioinformatics tools. It also holds significant potential for scaling to larger genomic datasets. This work validates EDVQE’s scalability and efficiency for genomic data while underscoring its utility in addressing complex genomics challenges, bridging quantum computing with practical bioinformatics workflows.
| phasing Completeness (%) | Switch error rate (%) | Hamming error rate (%) | |
| EDVQE | 100 | 0 | 0 |
| WhatsHap | 100 | 0 | 0 |
@@ 6 Acknowledgements
This work was supported by Shenzhen Science and Technology Program, China (JCYJ20241202123906009).

### v2

@@ Abstract
MaxCut is a canonical NP-hard combinatorial optimization problem in graph theory with broad applications ranging from physics to bioinformatics. Although variational quantum algorithms offer promising new approaches that may eventually outperform classical schemes, they suffer from resource constraints and trainability issues such as barren plateaus, making large-scale instances intractable on noisy intermediate-scale quantum devices. In this paper, we propose an enhanced distributed variational quantum eigensolver for large-scale MaxCut problems, which extends our prior distributed variational quantum eigensolver framework by integrating a novel hybrid classical-quantum perturbation strategy, enhances optimization scalability and efficiency. Our algorithm solves weighted MaxCut instances with up to 1000 vertices using only 10 qubits, and numerical results indicate that it consistently outperforms the Goemans-Williamson algorithm. We further employ a warm-start initialization strategy, seeding the algorithm with high-quality solutions from the Goemans-Williamson algorithm, with results confirming that the optimal classical solution can be effectively further improved. The practical utility of the proposed algorithm is further validated through its application to haplotype phasing on genome sequencing data of the human ABCA1 gene, producing high-quality haplotypes that rival those obtained by the Goemans-Williamson algorithm with projections. These results establish the proposed algorithm as a scalable, NISQ-compatible framework for near-term quantum-enhanced large-scale combinatorial optimization.
@@ 1 Introduction
For large-scale datasets, MaxCut becomes computationally intractable for classical heuristics, despite their effectiveness on smaller instances. This fundamental barrier drives the search for new computational paradigms, particularly quantum computing, which holds promise for outperforming classical methods on such hard problems. In the current noisy intermediate-scale quantum (NISQ) era[10], variational quantum algorithms (VQAs) [11, 12, 13, 14] represent the leading hybrid quantum-classical framework for tackling combinatorial optimization tasks. By employing a parameterized quantum circuit, also named ansatz, whose parameters are iteratively optimized by a classical optimizer to minimize or maximize a problem-encoded cost function, VQAs strike an attractive balance between expressiveness and hardware compatibility. Such as the variational quantum eigensolver (VQE)[15], originally developed for finding molecular ground states but readily adapted to Ising-type objectives, and the quantum approximate optimization algorithm (QAOA) [16], specifically designed for discrete combinatorial problems such as MaxCut. Both approaches are relatively resilient to noise when operated at shallow circuit depths, making them among the most viable candidates for near-term quantum advantage.
@@
However, such distributed frameworks are also plagued by the barren plateaus, a pervasive challenge in variational quantum algorithms, where the cost function’s gradient vanishes exponentially as the number of qubits increases [24, 25, 26, 27]. This decay severely impedes the classical optimization loop, often trapping it in suboptimal local minima and limiting scalability. The issues of vanishing gradients and local optima are likewise prevalent in classical optimization. In metaheuristic algorithms, a common and effective countermeasure is the deliberate introduction of perturbations. For instance, simulated annealing employs temperature-controlled stochastic acceptance [28], genetic algorithms utilize mutation operators [29], and variable neighborhood search relies on systematic neighborhood shifts [30, 31]. These strategies have all proven effective on challenging problems such as MaxCut and the traveling salesman problem (TSP) [32, 33]. In variational quantum algorithms, an analogous and highly effective perturbation-like technique is warm-starting. It leverages a high-quality classical solution to prepare a biased initial state, thereby accelerating convergence, facilitating escape from local minima, and increasing robustness against barren plateaus. Numerical experiments on small-scale graph instances further confirm that the warm-start method consistently outperforms the standard Goemans-Williamson algorithm[34, 35].
@@
A hybrid classical-quantum perturbation mesethod rooted in the MaxCut problem is proposed and integrated with the DVQE algorithm, enabling the solution of large-scale instances using merely 10 qubits, which significantly reduces hardware requirements while improving algorithmic performance.
@@
Simulation results reveal that the warm-start methodology consistently outperforms the Goemans-Williamson algorithm.
@@ 2.2 Haplotype phasing
Traditionally, haplotype information has been inferred statistically from population genotype data using linkage disequilibrium, as implemented in algorithms like PHASE[36]. However, the switch error rate is relatively high (approximately 5.4%) in unrelated individuals[37]. With the advancement of high-throughput sequencing technologies, particularly the emergence of long fragments and paired-end sequencing, direct haplotype phasing from sequencing data of a single individual has become feasible. The HapCUT algorithm [38] proposed by Bansal and Bafna (2008) represents an efficient and accurate combinatorial optimization method that transforms the problem into iteratively solving MaxCut on associated graphs, minimizing the minimum error correction (MEC) score. On the HuRef personal genome data, it achieved assembly accuracy significantly superior to greedy heuristics (20-25% lower MEC scores) and previous methods, reducing the switch error rate to 1.1-1.4%. This advancement provides more reliable complete haplotype information for personalized genomics, disease association studies, and precision medicine.
@@ 3.1 Framework of Enhanced Distributed Variational Quantum Eigensolver
Classical Neighborhood Search: Starting from , apply a classical neighborhood search algorithm (e.g., one based on single-vertex flips) to perform a local exploration, yielding an improved solution .
@@ 3.2 Distributed Variational Quantum Eigensolver
In practice, the Hamiltonian for binary optimization problems is typically expressed as a linear combination of terms:
@@
Moreover, each state of the subsystem is prepared and optimized on smaller quantum processors, while the global energy expectation is computed through the factorized expression.
@@ 3.3 Hybrid Classical-Quantum Perturbation
(9) where yields the identity operation , and results in , performing a SWAP operation up to a global phase. The parameter for each gate is initialized from a uniform distribution to ensure the system starts from a state that is only slightly perturbed from the CNS-1 solution. The subsequent variational optimization of these parameters enables probabilistic swapping of partition assignments, thereby facilitating exploration of more diverse solution spaces.
@@ 4.2 Performance on Weighted Complete Graphs
The performance of both algorithms, in terms of the normalized average cut value , is summarized in Figure 2 across a range of vertex counts. The results clearly demonstrate the superiority of the EDVQE algorithm over the standard GW algorithm under various projection configurations. Specifically, for problem sizes ranging from 100 to 1000 vertices, the EDVQE without hybrid classical-quantum perturbation, which represents the initial solution, consistently outperforms the standard GW algorithm employing a single projection. This observation demonstrates the inherent advantage of the base DVQE framework. Moreover, integrating hybrid classical-quantum perturbation within the EDVQE framework yields supplementary performance enhancements. To establish a comprehensive benchmark, we enhance the GW algorithm by increasing its number of projections, which correspondingly elevated its solution quality. As a result, our EDVQE approach maintains consistent superiority over the GW variant with 10 projections and, notably, outperforms the GW algorithm with 100 projections across instances ranging from 600 to 1000 vertices.
To further analyze the performance gains within the EDVQE framework, we break down the contributions of its three key components, which include the initial solution, the classical single-vertex neighborhood search (CNS-1), and the quantum two-vertex perturbation (QP-2). Table 1 summarizes the normalized average cut values of the initial solution, along with the absolute improvements from CNS-1 (CNS-1) and QP-2 (QP-2), together with their corresponding percentage enhancements. The results indicate that CNS-1 provides a consistent yet modest absolute improvement, with an average of 0.0025 (a relative gain of 0.089%). In contrast, QP-2 contributes a more substantial enhancement, delivering an average of 0.0043 (a relative gain of 0.150%)—approximately double that of the classical component. These findings demonstrate that both post-processing modules consistently improve upon the initial solution, with the quantum-based perturbation yielding notably greater gains.
| # of Nodes | Value | % Improvement | ||||
@@ 4.3 Performance on Weighted Cluster Graphs
Table 2 presents a component-wise analysis of the performance improvements in the EDVQE. The CNS-1 yields minimal improvement, with an absolute gain of 0.0050, corresponding to a relative gain of 0.085%. In contrast, the QP-2 contributes a substantially larger enhancement, attaining an absolute of 0.1386 and a relative gain of 2.168% relatively, exceeding the classical component by more than 25 times. This pronounced performance gap underscores the quantum module’s distinctive capability to identify cross-community optimal configurations via coordinated multi-qubit operations, thereby effectively addressing the limitations of classical local search in modular graph topologies.
| # of Nodes | Value | % Improvement | ||||
@@ 4.5 Performance for Haplotype phasing
| # of Nodes | GW () | EDVQE | |||
| Value | Best Value | Average Value | Success Rate | Average Approximation Ratio | |
@@
To assess phasing accuracy, the performance of EDVQE is systematically compared against the gold-standard small variant benchmark dataset derived from the Genome in a Bottle Consortium. As shown in Table 4, EDVQE achieves 100% phasing completeness with zero switch error rate and Hamming error rate, with independent validation using the bioinformatics tool WhatsHap confirming identical results. The comparative analysis demonstrates that EDVQE is viable for real-world haplotyping applications, accurately phasing heterozygous SNP variants into fully resolved and error-free haplotypes comparable to established bioinformatics tools. It also holds significant potential for scaling to larger genomic datasets. This work validates EDVQE’s scalability and efficiency for genomic data while underscoring its utility in addressing complex genomics challenges, bridging quantum computing with practical bioinformatics workflows.
| Phasing Completeness | Switch Error Rate | Hamming Error Rate | |
| EDVQE | 100% | 0 | 0 |
| WhatsHap | 100% | 0 | 0 |
@@ 6 Acknowledgements
This work was supported by Shenzhen Science and Technology Program, China (JCYJ20241202123906009), Guangdong Provincial Quantum Science Strategic Initiative (Grant No. GDZX2503001).